OpenStudy (anonymous):

If 6t+7a/4=t, then what is 4t in terms of a?

OpenStudy (anonymous):

6t+7a= 4t 7a=-2t (-2)7a=(-2)-2t 4t=-14a

OpenStudy (anonymous):

you can either solve for t in terms of a and then times 4 or just get 4t directly. 6t -t = -7*a/4 5t = -7*a/4 5*4/5* t = -7*a/4 *4/5 you can multiple a constant on both sides, it does not affect your answer, you get 4t = -7/5*a, and thats your answer, 4t = -7/5*a